OpenStudy (anonymous):

will Medal What is the solution to the system of equations? x=4y+6 x=3y=4 What is the solution to the system of equations? x+y=2 x-y+4

OpenStudy (studygurl14):

Hi :)

OpenStudy (studygurl14):

Is this a mistype? "x=3y=4"

OpenStudy (studygurl14):

and this too? "x-y+4"

OpenStudy (anonymous):

no sorry thats what it says

OpenStudy (anonymous):

@chosenmatt

OpenStudy (studygurl14):

oh, ok...well...then. that doesn't make any sense and it is unsolvable...

OpenStudy (anonymous):

wait im sorry it was a type-o its x-3y+4

OpenStudy (anonymous):

and its x+y=4

OpenStudy (anonymous):

@StudyGurl14 @skullpatrol

OpenStudy (studygurl14):

"x-3y+4" Isn't possible either. But x + y = 4 is, so i can help with that one

OpenStudy (studygurl14):

x + y = 2 x + y = 4 Are these together?

OpenStudy (anonymous):

yes

OpenStudy (studygurl14):

Then... 2 = 4 But... \(\large 2\neq 4\) So...there are no possible solutions. :)

OpenStudy (anonymous):

thank you

OpenStudy (studygurl14):

(I've gotten that answer on my stuff too, so don't worry that it is wrong)

OpenStudy (studygurl14):

yep :)